Please ignore the circled choices. I am double checking my work.
53 59 6 21 83 95 10 -1.05 1.0 2.0 2.5 40507599 percentile rank 20 (Q1) 5 Whats the missing raw score? a) 70 c)77 d 95 e) 98 6. Whats the missing z-score? -1.5 b) -1.0 ).5de) Whats the difference between the two missing percentile a) 0 b)S c) 10 d) 1se25 8. Since the mean is to the of the this distribution is skewed a) left, positively b) left, negatively righ, positively d) right, negatively e) none of these
Answer #1

5)

X = mean + z* sd

83 = mean + 1 * sd

95 = mean + 2 * sd

solving mean = 71

sd = 12

z = 0

X = mean = 71

option B)

6)

z = (X - mean)/sd = (53 - 71)/12

=   -1.5

7)

percentile = 100 * P(-1.5 < Z < -0.5)

= 25 (approx)

8)

a distribution that is skewed to the left , it has a mean smaller than its median.

mean = 71

median = 83

mean is left of median

option b) left, negative